Solve (1+yz)dx+x(z-x)dy-(1+xy)dz=0
1
Expert's answer
2020-08-25T12:55:05-0400

Let "z-x=u", then "z=x+u" and "dz=dx+du".

Rewrite the equation replacing "z" by "x+u":

"(1+xy+yu)dx+xudy-(1+xy)dx-(1+xy)du=0"

After simplification we obtain "yudx+xudy-(1+xy)du=0"

Note that "xdy+ydx=d(xy+1)", so we have "ud(1+xy)-(1+xy)du=0".

After dividing it by "(1+xy)^2" we obtain "d\left(\frac{u}{1+xy}\right)=-\frac{ud(1+xy)-(1+xy)du}{(1+xy)^2}=0", so "\frac{u}{1+xy}=C"

Since "u=z-x", we have "\frac{z-x}{1+xy}=C". After expressing "z" we obtain "z=x+C(1+xy)"

Answer: "z=x+C(1+xy)"